OpenStudy (anonymous):

How do you solve Lauren Morse works 7.5 hours per day. She spent a recent day making telephone calls, writing emails, and attending meetings. She spen twice as much time attending meetings as making telephone calls and 0.5 hour longer writing emails than making telephone calls. How many hours I'd she spend on each task?

OpenStudy (valpey):

7.5 = tc + we + am (...works 7.5 hours per day. She spent a recent day making telephone calls, writing emails, and attending meetings.) am = 2*tc (She spent twice as much time attending meetings as making telephone calls) we = 0.5 + tc 7.5 = tc + (0.5 + tc) + (2*tc) = 0.5 + 4*t.c.

OpenStudy (anonymous):

I'm not sure I understand it